# Down with the Hierarchy: The ‘H’ in HNSW Stands for “Hubs”

### v2

@@ Abstract.
Driven by recent breakthrough advances in neural representation learning, approximate near-neighbor (ANN) search over vector embeddings has emerged as a critical computational workload. With the introduction of the seminal Hierarchical Navigable Small World (HNSW) algorithm, graph-based indexes have established themselves as the overwhelmingly dominant paradigm for efficient and scalable ANN search. As the name suggests, HNSW searches a layered hierarchical graph to quickly identify neighborhoods of similar points to a given query vector. But is this hierarchy even necessary? A rigorous experimental analysis to answer this question would provide valuable insights into the nature of algorithm design for ANN search and motivate directions for future work in this increasingly crucial domain. To that end, we conduct an extensive benchmarking study covering more large-scale datasets than prior investigations of this question. We ultimately find that a flat navigable small world graph graph retains all of the benefits of HNSW on high-dimensional datasets, with latency and recall performance essentially identical to the original algorithm but with less memory overhead. Furthermore, we go a step further and study why the hierarchy of HNSW provides no benefit in high dimensions, hypothesizing that navigable small world graphs contain a well-connected, frequently traversed “highway” of hub nodes that maintain the same purported function as the hierarchical layers. We present compelling empirical evidence that the Hub Highway Hypothesis holds for real datasets and investigate the mechanisms by which the highway forms. The implications of this hypothesis may also provide future research directions in developing enhancements to graph-based ANN search.
1. Introduction
@@
Near neighbor search is a fundamental problem in computational geometry that lies at the heart of countless practical applications. From industrial-scale recommendation (Feng et al., 2022) to retrieval-augmented generation (Lewis et al., 2020) and even to computational biology (Zhao et al., 2024), numerous data-intensive tasks utilize similarity search at some location in the stack. As a result, similarity indexes are very well-studied (Guo et al., 2019; Malkov and Yashunin, 2016; Johnson et al., 2017; Aguerrebere et al., 2023; Jayaram Subramanya et al., 2019) with multiple large-scale benchmarks and leaderboards to compare techniques (Aumüller et al., 2018; Simhadri et al., 2022).
@@
Historically, the state-of-the-art for near neighbor search involved constructing sophisticated tree-based data structures, such as $kd$-trees (Bentley, 1975) and cover trees (Beygelzimer et al., 2006), that guaranteed exact solutions while avoiding a brute-force examination of all points. However, the recent advent of large-scale neural representation learning, including large language models (LLMs), places a significant strain on these classical methods that were developed to target a much lower-dimensional search space. In response, the community has turned to approximate search methods. While alternative approximate indexing methods such as locality-sensitive hashing (Indyk and Motwani, 1998) and product quantization (Jegou et al., 2010), have garnered significant interest, graph-based approaches generally achieve the strongest performance on established ANN benchmarks (Aumüller et al., 2018; Simhadri et al., 2022). Introduced in 2016, the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2016), emerged as one of the first high-performance graph-based search indexes at scale and still enjoys immense popularity to this day with over 4300 Github stars [^1] https://github.com/nmslib/hnswlib.
@@
Despite the immense popularity of HNSW, however, the algorithm suffers from multiple scalability bottlenecks. For instance, the hierarchical layers introduce a hefty memory overhead. Moreover, Malkov and Yashunin (2016) also note in Section 6 of their paper that the hierarchy can reduce the overall system throughput in distributed settings when compared to a flat NSW graph.
While the high cost of this hierarchy in HNSW is traditionally justified by the low latency of graph-based indexes, we ask the question of whether the hierarchy is even necessary in the first place. A growing body of results suggest that the hierarchy may be a vestigial artifact left behind by the lower-dimensional problems of the past. For example, Lin and Zhao (2019) published a set of experiments suggesting that the hierarchical component of HNSW is only beneficial for low-dimensional inputs ($d<32$). Other recent works have observed similar problems with the hierarchy. For example, Coleman et al. (2022) present an ablation study showing the same behavior in the appendix of their paper. This behavior seems to be an increasingly well-known “folklore” of the similarity search community, but there is currently an absence of a thorough investigation and analysis of this phenomenon.
An exhaustive benchmark comparison on the utility of the HNSW hierarchy would go a long way towards improving our understanding of graph-based similarity search in high-dimensional space and inform future research directions in developing improvements. However, there are several key difficulties with executing such a benchmarking study. The first issue is that running a meaningful ablation study requires first engineering a similarity search implementation over a flat NSW graph that reaches performance parity with HNSW and other state of the art approaches (Douze et al., 2024; Sun et al., 2024; Jayaram Subramanya et al., 2019). This is an increasingly nontrivial task, as the community has invested significant performance engineering efforts into these codebases such that a non-expert implementation has little hope of being competitive.
The second challenge is designing benchmarking experiments in a manner that avoids the confounding effect of the performance of a particular software implementation when making conclusions about the efficacy of the algorithm. Finally, benchmarking near-neighbor search can be highly nuanced. For example, it is possible that hierarchical structures are strongly beneficial for isolated slices of the input. These may sharply reduce the 99th percentile latency while not affecting the median.
@@
Perhaps most importantly, we still have no satisfactory understanding of why hierarchy does not help. Hierarchical structures are a mainstay of algorithm design, where a common trick is to reduce an $O(n)$ search process to a sublinear one by traversing a (balanced) hierarchy (Pugh, 1990; Guibas and Sedgewick, 1978; Mikolov et al., 2013; Cormen et al., 2022). Arguably, it is counterintuitive for this idea to fail to hold in the context of high-dimensional similarity search – especially when we have strong positive results that hierarchy helps in low dimensions (Beygelzimer et al., 2006; Dolatshah et al., 2015; Ram and Sinha, 2019; Lin and Zhao, 2019). Thus, an exhaustive benchmark and deeper analysis into the necessity of the hierarchy in HNSW would shed further light on the nature of algorithm design in high-dimensional spaces and thus may be of independent interest to the community as well.
@@
Figure 1. We hypothesize that in high dimensions, graph-based ANN indexes naturally form a “highway-feeder” structure, where a small subset of nodes and edges are easily reached, well-connected, and heavily traversed.
1.1. Contributions
@@
In this paper, we address the question of whether the hierarchical component of HNSW is truly necessary. Our central research question is, “Can we achieve the same performance on large-scale benchmarks with simply a flat navigable small world graph?” In summary, we make the following contributions:
@@
Why does hierarchy not help? We hypothesize that the hierarchy benefits decrease in high-dimensions due to hubness. Hubness is a high-dimensional phenomenon that causes a skewed distribution in the near-neighbor lists of search queries (Radovanovic et al., 2010). We hypothesize that hubness leads to preferential attachment in the similarity search graph, inducing the formation of easily-traversed highways that connect disparate regions of the graph. This hypothesis, which we call the Hub Highway Hypothesis, explains why we no longer need the hierarchy in high dimensions; we can simply traverse the intrinsic highway structure that naturally forms in high-dimensional spaces. To that end, we conduct a series of experiments to investigate the hub-highway hypothesis. Our results ultimately support this hypothesis, showing that hubness is responsible for driving the connectivity of similarity search graphs. This insight opens up exciting new research directions in graph construction, link pruning, and graph traversal.
@@
  • We release an implementation for a flattened version of HNSW [^2] https://github.com/nmslib/hnswlib, called FlatNav [^3] https://github.com/BlaiseMuhirwa/flatnav, that reaches performance parity with the original version with considerable memory savings.
@@
  • We demonstrate that hierarchy does not improve performance in either the median or tail latency case by building HNSW and FlatNav indexes over 13 popular benchmark datasets ranging in size from 1 million to 100 million vectors.
@@
  • We conduct an analysis of hubness phenomena in high-dimensional metric spaces and the resulting HNSW graphs, finding strong empirical support for the hub-highway hypothesis.
@@
Practical implications: Our benchmarks reveal that HNSW can be significantly optimized for modern high-dimensional embedding workloads. For instance, our implementation saves roughly $38\%$ and $39\%$ of peak memory consumption on two Big-ANN benchmark datasets compared to hnswlib (and sizable further headroom is likely). Our results confirm the folklore of the similarity search community, conclusively demonstrating that we can remove the hierarchy on high-dimensional inputs with impunity.
2. Background: Similarity Search & HNSW

### v3

@@ Abstract
Driven by recent breakthrough advances in neural representation learning, approximate near-neighbor (ANN) search over vector embeddings has emerged as a critical computational workload. With the introduction of the seminal Hierarchical Navigable Small World (HNSW) algorithm, graph-based indexes have established themselves as the overwhelmingly dominant paradigm for efficient and scalable ANN search. As the name suggests, HNSW searches a layered hierarchical graph to quickly identify neighborhoods of similar points to a given query vector. But is this hierarchy even necessary? A rigorous experimental analysis to answer this question would provide valuable insights into the nature of algorithm design for ANN search and motivate directions for future work in this increasingly crucial domain. We conduct an extensive benchmarking study covering more large-scale datasets than prior investigations of this question. We ultimately find that a flat navigable small world graph graph retains all of the benefits of HNSW on high-dimensional datasets, with latency and recall performance essentially identical to the original algorithm but with less memory overhead. Furthermore, we go a step further and study why the hierarchy of HNSW provides no benefit in high dimensions, hypothesizing that navigable small world graphs contain a well-connected, frequently traversed “highway” of hub nodes that maintain the same purported function as the hierarchical layers. We present compelling empirical evidence that the Hub Highway Hypothesis holds for real datasets and investigate the mechanisms by which the highway forms. The implications of this hypothesis may also provide future research directions in developing enhancements to graph-based ANN search.
Vector Databases, Approximate Near-Neighbor Search, HNSW
1 Introduction
@@
Near neighbor search is a fundamental problem in computational geometry that lies at the heart of countless practical applications. From industrial-scale recommendation (Feng et al., 2022) to retrieval-augmented generation (Lewis et al., 2020) and even to computational biology (Zhao et al., 2024), numerous data-intensive tasks utilize similarity search at some location in the stack. As a result, similarity indexes are very well-studied (Guo et al., 2019; Malkov & Yashunin, 2016; Johnson et al., 2017; Aguerrebere et al., 2023; Jayaram Subramanya et al., 2019) with multiple large-scale benchmarks and leaderboards to compare techniques (Aumüller et al., 2018; Simhadri et al., 2022).
@@
Historically, the state-of-the-art for near neighbor search involved constructing sophisticated tree-based data structures, such as $kd$-trees (Bentley, 1975) and cover trees (Beygelzimer et al., 2006), that guaranteed exact solutions while avoiding a brute-force examination of all points. However, the recent advent of large-scale neural representation learning, including large language models (LLMs), places a significant strain on these classical methods that were developed to target a much lower-dimensional search space. In response, the community has turned to approximate search methods. While alternative approximate indexing methods such as locality-sensitive hashing (Indyk & Motwani, 1998) and product quantization (Jegou et al., 2010), have garnered significant interest, graph-based approaches generally achieve the strongest performance on established ANN benchmarks (Aumüller et al., 2018; Simhadri et al., 2022). Introduced in 2016, the Hierarchical Navigable Small World (HNSW) algorithm (Malkov & Yashunin, 2016), emerged as one of the first high-performance graph-based search indexes at scale and still enjoys immense popularity to this day with over 4300 Github stars [^1] https://github.com/nmslib/hnswlib and deployments in major commercial search systems such as Apache Lucene (Xian et al., 2024) and Pinterest (Pinterest, 2021).
@@
Despite its popularity, HNSW has notable scalability issues; the hierarchical structure adds significant memory overhead and, as noted in (Malkov & Yashunin, 2016), can reduce throughput in distributed settings compared to flat NSW graphs. Although this overhead is often justified by the latency benefits of graph-based indexes, recent work questions whether the hierarchy is still necessary. (Lin & Zhao, 2019) report that the hierarchy only improves performance for low-dimensional data ($d<32$), and (Coleman et al., 2022) reach similar conclusions in their ablation study. These observations highlight the need for a more rigorous investigation into whether the hierarchy remains useful in modern, high-dimensional workloads.
@@
Perhaps most importantly, we still have no satisfactory understanding of why hierarchy does not help. Hierarchical structures are a mainstay of algorithm design, where a common trick is to reduce an $O(n)$ search process to a sublinear one by traversing a (balanced) hierarchy (Pugh, 1990; Guibas & Sedgewick, 1978; Mikolov et al., 2013; Cormen et al., 2022). Arguably, it is counterintuitive for this idea to fail to hold in the context of high-dimensional similarity search – especially when we have strong positive results that hierarchy helps in low dimensions (Beygelzimer et al., 2006; Dolatshah et al., 2015; Ram & Sinha, 2019; Lin & Zhao, 2019). Thus, an exhaustive benchmark and deeper analysis into the necessity of the hierarchy in HNSW would shed further light on the nature of algorithm design in high-dimensional spaces and thus may be of independent interest to the community as well.
@@
Figure 1: We hypothesize that, in high dimensions, graph-based ANN indexes naturally form a “highway-feeder” structure, where a small subset of nodes and edges are easily reached, well-connected, and heavily traversed.
1.1 Contributions
@@
In this paper, we study whether the hierarchical component of HNSW is truly necessary. Our central research question is, “Can we achieve the same performance on large-scale benchmarks with simply a flat navigable small world graph?” To that end, we organize the paper into two parts:
@@
Why does hierarchy not help? We hypothesize that the hierarchy benefits decrease in high-dimensions due to hubness. Hubness is a high-dimensional phenomenon that causes a skewed distribution in the near-neighbor lists of search queries (Radovanovic et al., 2010). We hypothesize that hubness leads to preferential attachment in the similarity search graph, inducing the formation of easily-traversed highways that connect disparate regions of the graph. This hypothesis, which we call the Hub Highway Hypothesis, explains why we no longer need the hierarchy in high dimensions; we can simply traverse the intrinsic highway structure that naturally forms in high-dimensional spaces. Our results ultimately show that hubness is responsible for driving the connectivity of similarity search graphs. This insight opens up exciting new research directions in graph construction, link pruning, and graph traversal.
@@
  • We release an implementation for a flattened version of HNSW, called FlatNav [^2] https://github.com/BlaiseMuhirwa/flatnav, that reaches performance parity with the original hierarchical version with considerable memory savings. To our knowledge, flatnav is the only actively maintained, high-performance NSW search library in the open-source ecosystem and thus fills a crucial void in the similarity search community.
@@
  • We demonstrate that hierarchy does not improve performance in either the median or tail latency case by building HNSW and FlatNav indexes for 13 popular benchmark datasets ranging in size from 1 million to 100 million vectors.
@@
  • We present strong scientific evidence for the hub-highway hypothesis, drawing empirical support from analysis of hubness phenomena in high-dimensional metric spaces and the resulting HNSW graphs.
@@
Practical implications: Our benchmarks reveal that HNSW can be significantly optimized for modern high-dimensional embedding workloads. For instance, as we show in Table 6 in the appendix, our implementation saves roughly $38\%$ and $39\%$ of peak memory consumption during index construction on two Big-ANN benchmark datasets compared to hnswlib (and sizable further headroom is likely). Our results confirm the folklore of the similarity search community, conclusively demonstrating that we can remove the hierarchy on high-dimensional inputs with impunity.
2 Related Work
2.1 Near-Neighbor Benchmarks
Recent years have seen the emergence of large-scale benchmarks for the $k$-NNS problem. ANN Benchmarks (Aumüller et al., 2018) established the first standard evaluation framework, expanding over time to cover 30+ methods across 9 datasets. However, the ANN Benchmark datasets are relatively small, with around one million points. To better reflect real-world scale, Big ANN Benchmarks (Simhadri et al., 2022) introduced five billion-scale datasets. Nevertheless, these benchmarks emphasize overall throughput and omit tail latency metrics like the 99th percentile, which are essential for evaluating hierarchical components such as those in HNSW.
Hierarchy Studies: Many top-performing graph-based ANN algorithms, including HNSW (Malkov & Yashunin, 2016), ONNG (Iwasaki & Miyazaki, 2018), PANNG (Iwasaki, 2016), and HCNNG (Munoz et al., 2019), rely on hierarchical structures. However, recent work has questioned this design. (Dobson et al., 2023) show that HNSW can underperform both HCNNG (with a shallower hierarchy) and DiskANN (Jayaram Subramanya et al., 2019) (which lacks one). (Lin & Zhao, 2019) find hierarchy helpful only in low-dimensional synthetic settings ($d<32$), but study few real-world datasets and do not explain the observed failure modes. We aim to address these gaps by reproducing prior results (Malkov & Yashunin, 2016; Lin & Zhao, 2019) with our own implementation and extending the analysis to larger datasets with a focus on when and why hierarchy matters.
2.2 Hubness in High Dimensional Spaces
Astute readers might observe that the HNSW graph construction algorithm does not explicitly enforce the small world property and instead adds edges between nodes based on their proximity in the metric space. The connection between proximity and the small world property arises due to hubness.
Hubness is a property of high-dimensional metric spaces where a small subset of points (the “hubs”) occur a disproportionate number of times in the near-neighbor lists of other points in the dataset (Radovanovic et al., 2010). In other words, a small fraction of nodes are highly connected to other points in the near-neighbor graph. The concentration of distance and measure in high-dimensional spaces provide good intuition for how hubness can arise in a datasets. The concentration of distances is a well-studied phenomenon where the expected $\ell_{2}$ distance between independent and identically distributed (i.i.d) vectors grows with $\sqrt{d}$ while the variance tends to a constant as $d$ approaches infinity (Talagrand, 1994). As a result, the $\ell_{2}$ distance loses its discriminative power as $d$ increases, a fact which also holds for $\ell_{p}$ and fractional norms (François et al., 2007). Concentration of measure is a high-dimensional property where random distributions have most of their mass near the boundary of their domain. Taken together, these facts suggest that hubs will form at extrema of high-dimensional datasets - a result which holds true empirically (Low et al., 2013).
Due to undesirable consequences of the hubness phenomenon, such as poor clustering quality, a large body of work has focused on hubness reduction strategies. For instance, (Zelnik-Manor & Perona, 2004) introduced local scaling which scales distances $d(\mathbb{x},\mathbb{y})$ by accounting for local neighborhood information. Interestingly, our work stands in contrast to this literature on hubness reduction by presenting a case study where hubs provide tangible value in an algorithmic setting, namely in accelerating greedy traversal in near neighbor proximity graphs. This result may be of independent interest to machine learning and algorithms researchers as well.
3 FlatNav Benchmarking Experiments
In this section, we report the results of our benchmarking study comparing the performance of flat HNSW search to hierarchical search on a suite of standard high-dimensional benchmark datasets drawn from real machine learning models. We fix the implementation in our experimental design such that the same code is used to construct the indexes. In particular, we use the hnswlib library as our baseline HNSW implementation. To benchmark the flat NSW index performance, we extract the bottom layer from hnswlib after constructing the full hierarchical graph and reimplement HNSW’s greedy search heuristic over the flat graph via flatnav.
One natural question that arises from this experimental design is whether the hierarchical component of HNSW might still be useful during construction even if the base layer suffices for search. We find that there is no difference in performance between these settings. Specifically, we include additional results comparing HNSW to flat graphs constructed from scratch (without the hierarchy) in Appendix E where we see identical results to those reported in this section of the paper. In this section, we focus on extracting the base layer from the hnswlib graph to reduce any potential confounding effects from implementation differences. Nevertheless, we obtain identical results regardless of how the base graph is constructed.
3.1 Datasets and Compute
We utilize the benchmark datasets released through the popular leaderboards ANN Benchmarks (Aumüller et al., 2018) (MIT Licensed) and Big ANN Benchmarks (MIT Licensed) (Simhadri et al., 2022). The specific datasets and their associated statistics are presented in Table 4. For the Big ANN Benchmark datasets, we consider both the 10M and 100M collection of vectors, for which the ground truth near neighbors have previously been computed and released. We did not experiment with the largest Big ANN datasets with 1 billion vectors since constructing HNSW indexes at this scale requires over 1.5TB of RAM, which exceeded our compute resources. In this section, we include our benchmarking results for the four 100M-scale datasets available through Big ANN Benchmarks. We see that our flat HNSW implementation achieves performance parity with the hierarchical HNSW implementation.
3.2 Latency Results
3.2.1 BigANN Benchmarks (Simhadri et al., 2022)
In Figures 3 and 3, we compare latency metrics for HNSW and FlatNav at the 50th and 99th percentile for the four 100M datasets from BigANN benchmarks listed in Table 4. All of our results support the conclusion that flatnav achieves nearly identical performance to hnswlib.
From the results in Figures 3 and 3, we observe that there is no consistent and discernable gap between FlatNav and HNSW in both the median and tail latency cases. These results suggest that the hierarchical structure of HNSW provides no tangible benefit on practical high-dimensional embedding datasets.
Figure 2: p50 Latency vs. Recall. FlatNav performs nearly identically to HNSW.
Figure 3: p99 Latency vs. Recall. FlatNav performs nearly identically to HNSW.
3.2.2 ANN Benchmarks (Aumüller et al., 2018)
We repeat the same experimental setup comparing HNSW and FlatNav on the ANN Benchmark datasets listed in Table 4. In Figures 5 and 5, we report the p50 and p99 latency of all of the non-GloVe ANN Benchmarks. Although these datasets are smaller in scale than the BigANN Benchmarks, we still see no discernible difference in latency between HNSW and FlatNav which supports our hypothesis that the vector dimensionality and not the size of the collection is the main driver of eliminating the need for hierarchical search in small world graphs. We see further evidence of this idea in Figure 6, which confirms that there is no clear performance benefit provided by the hierarchy of HNSW.
Figure 4: p50 Latency vs. Recall. FlatNav performs nearly identically to HNSW.
Figure 5: p99 Latency vs. Recall. FlatNav performs nearly identically to HNSW.
Figure 6: p50 and p99 Latency vs. Recall for HNSW and FlatNav over GloVe datasets.
Figure 7: Log-normalized Node access count distribution $P_{m}(\mathbb{x}_{i})$ for datasets using angular (left) and $l_{2}$ (right) distances.
4 The Hub Highway Hypothesis
We now turn our attention to studying why the hierarchy appears to provide no benefit in the search process. In our experiments, we observed that a small fraction of nodes appear in the set of near neighbors for a disproportionate number of other vectors. We thus conjecture that the hub structure prevalent in high-dimensional data performs the same functional role as the hierarchy.
Hypothesis (Hub Highways).
4.1 Methodology
  1. 1.
    Some nodes are visited by queries much more frequently than others. The relative popularity of these hub nodes is explained by the hubness phenomenon that arises in high dimensions.
  2. 2.
    The hub nodes form a well-connected subgraph of hubs (the highway network).
  3. 3.
Empirical measures of hubness: To support the first part of our argument, we require a formal characterization of hubness. Following (Radovanovic et al., 2010), let $\mathbb{x},\mathbb{x}_{1},\ldots,\mathbb{x}_{n}$ be vectors drawn from the same probability distribution supported on $\mathcal{S}\subseteq\mathbb{R}^{d}$, and let $\phi:\mathcal{S}\times\mathcal{S}\to\mathbb{R}$ be a distance function. For $1\leq i,k\leq n$, define $p_{i,k}(\mathbb{x})=1$ if $\mathbb{x}$ is among the $k$-NN set of $\mathbb{x}_{i}$ under $\phi$, and 00 otherwise. Define $N(\mathbb{x})\coloneqq\sum_{i=1}^{n}p_{i,k}(\mathbb{x})$, the number of vectors $\mathbb{x}_{i}$ for which $\mathbb{x}$ appears in their $k$-NN set.
Given a dataset $\mathcal{D}$, the values $\{N(\mathbb{x})\}_{\mathbb{x}\in\mathcal{D}}$ form a distribution $N_{k}$, whose skewness is given by $S_{N_{k}}=\mathbb{E}[(N_{k}-\mu_{N_{k}})^{3}]\,/\,\sigma_{N_{k}}^{3}$. We use $S_{N_{k}}$ to measure the hubness of a dataset.
This measure characterizes the asymmetry of the $k$-occurrence distribution $N_{k}$, and it is the metric most often used to estimate the presence of hubs. The more skewed the distribution of $N_{k}$, the greater the chance that a small number of vectors (hubs) will occur in the $k$-NN sets of other vectors.
Synthetic and ANN Benchmark Datasets: We use real and synthetic datasets as shown in Table 5 to study the Hub-Highway Hypothesis. In addition to a subset of ANN Benchmark datasets, we generate synthetic datasets by drawing vectors from the standard normal distribution.
4.2 Skewness of the Node Access Distribution
The goal of this study is to support our claim that some nodes are visited far more frequently, and that this process is driven by hubs in the metric space. We examine the discrete distribution of the number of times each node in the index is visited during search given a fixed number of queries, which we write as $P_{m}(\mathbb{x}_{i})$ (i.e., node $\mathbb{x}_{i}$ is visited $m$ times). If $P_{m}(\mathbb{x}_{i})$ is right-skewed, it means that some nodes are very frequently visited.
Table 1: Similarity search index parameters
$m$ $ef$-construction $ef$-search $k$
32 100 200 100
Figure 7 shows the log-normalized node access count distribution for different datasets. We observe that as the dimension $d$ increases, this distribution becomes right-skewed for $\ell_{2}$ distance-based datasets. While this is strong evidence for the first part of our claim, it leaves open the possibility that some mechanism other than hubness is driving our observations. To control for this possibility, we also study the cosine distance, which is known to have anti-hub properties that prevent hub formation (Radovanovic et al., 2010). We find that the cosine distance does not have a dramatic skew, even for $d\in\{1024,1536\}$. The increased skewness of the $P_{m}(\mathbb{x}_{i})$ distribution as $d$ increases demonstrates that highway nodes become increasingly prevalent as the dimension increases, and the differences between the $\ell_{2}$ and cosine results suggest that the hubness phenomenon is responsible for the formation of the highway nodes.
4.3 Subgraph Connectivity of the Hub-Highway Nodes
In this section we present empirical evidence confirming that hub-highway nodes exhibit strong connectivity in the graph. We begin by explaining our procedure to identify hubs. Let $\{\mathbb{x}_{i}\}_{i=1}^{n}$ be the vectors in a similarity search index for a dataset $\mathcal{D}$.
[…]
  • We wish to estimate the likelihood that a randomly-chosen out-neighbor of a hub node is, itself, a hub. To do so, we examine the 1-hop out-degree expansion of each hub and count the number of adjacent hub nodes. This yields a discrete distribution for the number of hubs to which each hub node is connected.
We test whether hubs differ from non-hubs in connectivity behavior by using statistical tests on the two distributions. Under the null hypothesis, both groups are equally likely to connect to other hubs; the alternative asserts that hubs preferentially attach to hubs.
We apply both a two-sample $t$-test and the Mann-Whitney U-test (Mann & Whitney, 1947), the latter of which is more suitable for ANN datasets since it makes no normality assumption. With dataset sizes $n>10^{3}$, Table 3 reports results using the top $5\%$ of nodes as hubs. At a 0.05 significance threshold, we reject the null in all but five cases using the U-test (and all but six with the $t$-test). Effect sizes are largest in synthetic $\ell_{2}$ datasets, likely due to their stronger hubness under this metric.
Using a stricter top-$1\%$ threshold (Table 3), we reject the null in all but one case for both tests, with substantially larger effect sizes. These findings indicate that the most prominent hubs tend to form tightly connected subgraphs, consistent with our Hub Highway Hypothesis.
Table 2: Two-sample $t$-test and Mann-Whitney U-test results. Hub nodes are selected using the P95 threshold of the node access distribution.
Table 3: Two-sample $t$-test and Mann-Whitney U-test results. Hub nodes are selected using the P99 threshold of the node access distribution.
Dataset Dim Mann-Whitney Two-Sample $t$-Test Effect Size IID Normal (Angular) 16 0.0006 0.0006 0.1745 IID Normal (L2) 16 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.6621 IID Normal (Angular) 32 0.0347 0.0347 0.0972 IID Normal (L2) 32 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8173 IID Normal (Angular) 64 0.0359 0.0417 0.0927 IID Normal (L2) 64 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8725 IID Normal (Angular) 128 0.0093 0.0070 0.1316 IID Normal (L2) 128 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8428 IID Normal (Angular) 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.3110 IID Normal (L2) 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8582 IID Normal (Angular) 1024 0.1472 0.1318 0.0598 IID Normal (L2) 1024 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8314 IID Normal (Angular) 1536 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.2356 IID Normal (L2) 1536 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.8568 GloVe 100 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.7642 NYTimes 256 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.9305 GIST 960 $\!<\!10^{-5}$ $\!<\!10^{-5}$ 0.6829 Yandex-DEEP 96 0.0013 0.0013 0.1614 Microsoft-SpaceV 100 0.0011 0.0011 0.1644
4.4 Hub-Highway Nodes Enable Fast Traversal
Our final question is whether the highway nodes allow queries to quickly traverse the similarity search graph. While it is not surprising that a well-connected subgraph of frequently visited nodes would enable this behavior, it is not necessarily the case that queries would use the highway in the way predicted by our hypothesis, namely to quickly identify a neighborhood for deep exploration. To investigate this question, we track the sequence of nodes visited during beam search for several thousand queries. This allows us to determine the fraction of time spent within hub nodes in different phases of search.
Since beam search takes a variable number of steps for each query, we normalize by the total search length when presenting the results. More formally, suppose that $\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{l}$ is a length-$l$ sequence of such nodes visited by a query. We use the hub node assignment heuristic discussed in section 4.3 to label $h(\mathbb{x}_{i})$ each of these nodes as hubs / non-hubs. We then split the sequence into bins and compute the prevalence of hubs in each bin. For bin $B_{i}$, this is given by
$\left(\frac{1}{|B_{i}|}\right)\sum_{\mathbb{x}_{j}\in B_{i}}h(\mathbb{x}_{j})$
Figure 8 shows results for the Gist, GloVe, Microsoft SpaceV and Yandex-DEEP benchmark datasets. We observe that queries tend to concentrate in the highway structures early in search, shown by the high percentage of hub nodes visited in the first 5-10% of the search steps. This result suggests that the highway allows queries to quickly navigate the similarity search graph until they find the region of the graph best suited for deep exploration. The propensity of the query to visit hubs appears to be tied to the hubness properties of the dataset. For example, the GloVe dataset uses the angular distance, has less pronounced hubs (Figure 7), and queries spend a lower percentage of their time in hub nodes in this dataset (Figure 8(b)). On the other hand, the GIST dataset has some of the highest highway utilization rates and is also our highest-dimensional $\ell_{2}$ dataset.
(a) Gist
(b) GloVe
(c) Microsoft SpaceV
(d) Yandex Deep
Figure 8: Highway nodes allow queries to traverse the graph faster.
4.5 Discussion
Our sequence of experiments provide substantial evidence supporting the Hub Highway Hypothesis. Although it has long been established that the hubness phenomenon negatively affect common applications, such as clustering and even near neighbor search recall, existing ANNS methods, such as HNSW and NSG (Zhao et al., 2023) mostly emphasize algorithmic improvements and performance optimizations only. We have shown that leveraging the inherent structures in the data, particularly hub-highway occurrences, should be central to the design of new scalable similarity search indexes.
Initialization is a recurring and popular research direction for graph-based near-neighbor search, and leading algorithms vary greatly in their initialization techniques. The research question dates back to the seminal 1993 paper by Arya and Mount (Arya & Mount, 1993), which conjectured that clever search initialization – in their case, via $kd$-tree – could improve performance over random initialization. Over the following three decades, the research community has investigated diverse stratified sampling based on clusters (Sebastian & Kimia, 2002), vantage-point trees (Iwasaki, 2010), seeds formed from the graph expansion of $kd$-trees (Iwasaki & Miyazaki, 2018) and previously-visited nodes (Wang & Li, 2012), hierarchical graphs (Malkov & Yashunin, 2016), hierarchical clustering (Munoz et al., 2019), dataset medoids (Jayaram Subramanya et al., 2019), and several other methods before finally returning to cluster-stratified candidates (Oguri & Matsui, 2024; Ni et al., 2023) and random entry nodes (Jaiswal et al., 2022).
5 Conclusion
Approximate near neighbor search has become an increasingly crucial computational workload in recent years with the seminal Hierarchical Navigable Small World (HNSW) algorithm continuing to garner significant interest and adoption from practitioners. We present the first comprehensive study on the utility of the hierarchical component of HNSW over numerous large-scale datasets and performance metrics. Ultimately, we find that the hierarchy of HNSW provides no clear benefit on high-dimensional datasets and can be removed without any discernible loss in performance while providing memory savings.
While similar observations have been made before in the literature (Lin & Zhao, 2019; Coleman et al., 2022; Dobson et al., 2023), we are, to our knowledge, the first to conduct an exhaustive study over modern benchmark datasets and taking extensive care to compare implementations with performance engineering parity. Furthermore, we go beyond prior works and study why the hierarchy does not help, culminating in our introduction of the Hub Highway Hypothesis, an empirical result on how proximity graphs built over high-dimensional metric spaces leverage a small subset of well-connected nodes to traverse the network quickly. We believe that our results provide immediate implications for practitioners seeking to save memory or simplify their vector database implementations and we look forward to further partnering with the community on these endeavors.
Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under Grant No. 2313998. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation.
Appendix
Appendix A Background: Similarity Search & HNSW
A.1 Similarity Search